# Fano threefolds in positive characteristic III

### v3

@@ 1. Introduction
This article is the third part of our series of papers. In the first and second parts [TanI], [TanII], we studied Fano threefolds $X$ in positive characteristic with $\rho(X)=1$. In this paper, we classify primitive Fano threefolds $X$ in positive characteristic with $\rho(X)\geq 2$. More precisely, the main theorem is as follows.
@@
Let $X$ be a primitive Fano threefold with $\rho(X)\geq 2$. Then $X$ is isomorphic to one of the following varieties.
@@ 3-27 $3$ $48$ $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$
    $W$ is a smooth prime divisor on $\mathbb{P}^{2}\times\mathbb{P}^{2}$ of bidegree $\mathcal{O}(1,1)$. Note that such a threefold is unique up to isomorphisms (Lemma 5.16).
@@
Let $X$ be a Fano threefold with $\rho(X)=2$. Then $X$ is isomorphic to one of the following varieties.
@@ 2-3 $8$ blowup of $V_{2}$ along an elliptic curve of degree $2$ $D_{1}+E_{1}$
2-4 $10$ blowup of $\mathbb{P}^{3}$ along a curve of genus $10$ degree $9$ $D_{1}+E_{1}$
@@ 2-5 $12$ blowup of $V_{3}$ along an elliptic curve of degree $3$ $D_{1}+E_{1}$
2-7 $14$ blowup of $Q$ along a curve of genus $5$ degree $8$ $D_{1}+E_{1}$
@@ 2-14 $20$ blowup of $V_{5}$ along an elliptic curve of degree $5$ $D_{1}+E_{1}$
2-15 $22$ blowup of $\mathbb{P}^{3}$ along a curve of genus $3$ and degree $6$ $E_{1}+E_{3}\,{\rm or}\,E_{4}$
@@ 2-36 $62$ $\mathbb{P}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(2))$ $C_{2}+E_{5}$
(I) Conic bundles. In characteristic zero, the discriminant divisor $\Delta_{f}$ of a threefold conic bundle $f:X\to Y$ is reduced and normal crossing. However, both the conclusions fail in characteristic two. Because of this, we shall need some minor modifications of the proofs. For example, if $Y=\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\Delta_{f}\sim\mathcal{O}(2,0)$, then we immediately get a contradiction in characteristic zero, because $\Delta_{f}$ does not contain any smooth rational curve as a connected component.
@@
Acknowledgements: This paper is based on the master thesis [Asa] of the first author, in which the primitive case of odd characteristic is carried out. The second author was funded by JSPS KAKENHI Grant numbers JP22H01112 and JP23K03028.
@@ 2.2. Split double covers
    If $p\neq 2$, then the branched divisor $D$ satisfies $\mathcal{O}_{Y}(D)\simeq\mathcal{L}^{\otimes 2}$.
@@
    Both (a) and (b) hold.
    1. (a)
      $Y$ is $F$-split, i.e., $\mathcal{O}_{Y}\to F_{*}\mathcal{O}_{Y}$ splits as an $\mathcal{O}_{Y}$-module homomorphism, where $F:Y\to Y$ denotes the absolute Frobenius morphism.
    2. (b)
      $H^{1}(Y,\mathcal{O}_{Y}(E))=0$ for any effective Cartier divisor $E$ on $Y$.
@@
It is well known that $f$ is split when (1) holds. Note that $\mathcal{O}_{Y}\to f_{*}\mathcal{O}_{X}$ splits if and only if
@@
splits. Since the extension class corresponding to this exact sequence is contained in $H^{1}(Y,\mathcal{L})$, (2) implies that $f$ is split.
Assume (3). If $f$ is inseparable, then the absolute Frobenius morphism $F:Y\to Y$ of $Y$ factors through $f:X\to Y$:
@@
and hence the splitting of $\mathcal{O}_{Y}\to F_{*}\mathcal{O}_{Y}$ implies the splitting of $\mathcal{O}_{Y}\to f_{*}\mathcal{O}_{X}$. Therefore, we may assume that $f:X\to Y$ is separable. In this case, the trace map ${\operatorname{Tr}}:f_{*}\mathcal{O}_{X}\to\mathcal{O}_{Y}$ is nonzero, because the field-theoretic trace map ${\operatorname{Tr}}:K(X)\to K(Y)$ is given by ${\operatorname{Tr}}(b)=b+\sigma(b)$ for the Galois involution $K(X)\to K(X)$, and hence ${\operatorname{Tr}}(b)\neq 0$ for $b\in K(X)\setminus K(Y)$. As the composite $\mathcal{O}_{Y}$-module homomorphism $\mathcal{O}_{Y}\hookrightarrow f_{*}\mathcal{O}_{X}\xrightarrow{{\operatorname{Tr}}}\mathcal{O}_{Y}$ is zero, we get a nonzero $\mathcal{O}_{Y}$-module homomorphism $\mathcal{L}^{-1}\simeq f_{*}\mathcal{O}_{X}/\mathcal{O}_{Y}\to\mathcal{O}_{Y}$. Therefore, $H^{0}(Y,\mathcal{L})\neq 0$. In this case, (b) implies $H^{1}(Y,\mathcal{L})=0$, i.e., (2) holds. Therefore, (3) implies that $f$ is split. ∎
@@ 2.3. Brauer groups
Let $f:X\to Y$ be a projective morphism of smooth varieties. Fix $r\in\mathbb{Z}_{>0}$. Assume that ${\operatorname{Br}}(Y)=0$ and any fibre of $f$ is isomorphic to $\mathbb{P}^{r}$. Then there exists a vector bundle $E$ of rank $r+1$ such that $X$ is isomorphic to $\mathbb{P}(E)$ over $Y$.
@@
Fix a closed point $y\in Y$ and set $A:=\mathcal{O}_{Y,y}$. For the completion $\widehat{A}=\widehat{\mathcal{O}}_{Y,y}$, we have an isomorphism $X\times_{Y}{\operatorname{Spec}}\,\widehat{A}\simeq\mathbb{P}^{r}_{\widehat{A}}$ over $\widehat{A}$ [Ser06, Corollary 1.2.15]. By Artin’s approxximation theorem, We obtain an $A^{h}$-isomorphism $X\times_{Y}{\operatorname{Spec}}\,A^{h}\simeq\mathbb{P}^{r}_{A^{h}}$ over the henselisation $A^{h}=\mathcal{O}_{Y,y}^{h}$. Therefore, there exists an étale surjective morphism $Y^{\prime}\to Y$ such that $X\times_{Y}Y^{\prime}\simeq\mathbb{P}^{r}_{Y^{\prime}}$.
@@
This implies that the ${\rm PGL}_{r+1}$-torsor $f:X\to Y$ comes from ${\rm GL}_{r+1}$-torsor. Since any ${\rm GL}_{r+1}$-torsor is nothing but a projective space bundle $\mathbb{P}_{Y}(E)$, the assertion holds by the following facts.
@@
  • $H^{1}_{{\operatorname{\acute{e}t}}}(Y,{\rm PGL}_{r+1})$ consists of the isomorphism classes of $\pi:W\to Y$ such that $W$ is a flat proper morphism such that $W\times_{Y}Y^{\prime}\simeq P^{r+1}_{Y^{\prime}}$ for some étale surjective morphism $Y^{\prime}\to Y$/. $\mathbb{P}^{r+1}$.
@@ 3.1. Types of extremal rays
    For any curve $C$ on $X$, $[C]\in R$ if and only if $f(C)$ is a point.
@@
    Assume that $R$ is of type $C$. Then the following hold.
@@ 3.2. Del Pezzo fibrations
    For a closed point $y\in Y$, the effective Cartier divisor $f^{*}y$ is a prime divisor, i.e., any scheme-theoretic fibre of $f$ is geometrically integral.
@@
    $X_{\overline{K}}$ is a projective canonical del Pezzo surface, i.e., $X_{\overline{K}}$ is a projective normal surface such that $V_{\overline{K}}$ has at worst canonical singularities and $-K_{V_{\overline{K}}}$ is ample.
@@
    ${\operatorname{Pic}}X_{\overline{K}}$ is a finitely generated free abelian groups. Furthermore, ${\operatorname{Pic}}X_{K}\simeq\mathbb{Z}$.
@@
    If $K_{X_{K}}^{2}=9$, then $f$ is a $\mathbb{P}^{2}$-bundle, i.e., there exists a locally free $F$ on $Y$ of rank $3$ such that $X$ is isomorphic to $\mathbb{P}_{Y}(F)$ over $Y$. In particular, any fibre is $\mathbb{P}^{2}$.
@@
$\Delta(X_{\overline{K}},\mathcal{O}_{X}(D)|_{X_{\overline{K}}})=2+\mathcal{O}_{X}(D)^{2}-h^{0}(X_{\overline{K}},\mathcal{O}_{X}(D)|_{X_{\overline{K}}})=0.$
@@ 3.3. Contraction morphisms
The assertions (1) and (2) follow from [TanI, Theorem 2.4] (cf. [Kaw21, Corollary 3.7]). The assertion (3) holds by $\chi(X,\mathcal{O}_{X})=1$ and $\chi(X,\mathcal{O}_{X})=\frac{1}{24}(-K_{X}\cdot c_{2}(X))$.
@@
On the other hand, by the exact sequences
$0\longrightarrow\mathcal{O}_{X}(-D)\longrightarrow\mathcal{O}_{X}\longrightarrow\mathcal{O}_{D}\longrightarrow 0$
and
$0\longrightarrow\mathcal{O}_{X}\longrightarrow\mathcal{O}_{X}(D)\longrightarrow\mathcal{O}_{D}(D)\longrightarrow 0,$
we have
@@
    If $X$ is a Fano threefold, then $R^{i}f_{*}\mathcal{O}_{X}=0$ for any $i>0$. Furthermore, $H^{j}(X,\mathcal{O}_{X})\simeq H^{j}(Y,\mathcal{O}_{Y})$ for any $j\geq 0$.
@@
(1) If $\dim X_{y}\leq 1$ for any closed point $y\in Y$, then $R^{i}f_{*}\mathcal{O}_{X}=0$ holds by [Tan15, Theorem 0.5]. We may assume that $\dim X_{y}\geq 2$ for some closed point $y\in Y$. In particular, $\dim Y=0$ or $\dim Y=3$, as we assume $\dim Y\neq 1$. If $\dim Y=0$, then we obtain $R^{i}f_{*}\mathcal{O}_{X}=0$ by Lemma 3.8.
Assume that $\dim Y=3$. By $\dim X_{y}\geq 1$, $f(E)$ is a point for $E:={\operatorname{Ex}}(f)$. Then we have an exact sequence for any $r\in\mathbb{Z}_{\geq 0}$:
@@
(2) By (1), we may assume that $\dim Y=1$. Pick general closed point $Q_{1},...,Q_{n}$ on $Y$. Set $S_{1}:=f^{*}Q_{1},...,S_{n}:=f^{*}Q_{n}$, and $D_{Y}:=Q_{1}+\cdots+Q_{n}$. By the Serre vanishing theorem
@@
Since each $S_{i}$ is a canonical del Pezzo surface (Proposition 3.5(2)), we have $H^{i}(S_{\ell},\mathcal{O}_{S_{\ell}})=0$ for every $\ell$ and every $i>0$. It holds that $H^{i}(X,\mathcal{O}_{X})=0$ for any $i>0$, which implies $H^{i}(X,\mathcal{O}_{X}(f^{*}D_{Y}))=0$ for any $i>0$. ∎
@@
Set $\mathcal{O}_{D}(D):=\mathcal{O}_{X}(D)|_{D}$. We have the exact sequence
$0\longrightarrow\mathcal{O}_{X}(-D)\longrightarrow\mathcal{O}_{X}\longrightarrow\mathcal{O}_{D}\longrightarrow 0,$
which yields
$0\longrightarrow\mathcal{O}_{X}\longrightarrow\mathcal{O}_{X}(D)\longrightarrow\mathcal{O}_{D}(D)\longrightarrow 0$
Then the following hold:

### v4

@@ 1. Introduction
This article is the third part of our series of papers. In the first and second parts [TanI], [TanII], we studied Fano threefolds $X$ in positive characteristic with $\rho(X)=1$, where $\rho(X)$ denotes the Picard number of $X$. In this paper, we classify primitive Fano threefolds $X$ in positive characteristic with $\rho(X)\geq 2$ (for the definition of primitive Fano threefolds, see Subsection 2.1(6)). More precisely, the main theorem is as follows.
@@
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $X$ be a primitive Fano threefold over $k$ with $\rho(X)\geq 2$. Then $X$ is isomorphic to one of the following varieties.
@@ 3-27 $3$ $48$ $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$
    $W$ is a smooth prime divisor on $\mathbb{P}^{2}\times\mathbb{P}^{2}$ of bidegree $(1,1)$. Note that such a threefold is unique up to isomorphisms (Lemma 5.16).
@@
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $X$ be a Fano threefold over $k$ with $\rho(X)=2$. Then $X$ is isomorphic to one of the following varieties.
@@ 2-3 $8$ blowup of $V_{2}$ along an elliptic curve of degree $2$ $D_{1}+E_{1}$
2-4 $10$ blowup of $\mathbb{P}^{3}$ along a curve of genus $10$ and degree $9$ $D_{1}+E_{1}$
@@ 2-5 $12$ blowup of $V_{3}$ along an elliptic curve of degree $3$ $D_{1}+E_{1}$
2-7 $14$ blowup of $Q$ along a curve of genus $5$ and degree $8$ $D_{1}+E_{1}$
@@ 2-14 $20$ blowup of $V_{5}$ along an elliptic curve of degree $5$ $D_{1}+E_{1}$
[…]
@@ 2-36 $62$ $\mathbb{P}(\mathcal{O}_{\mathbb{P}^{2}}\oplus\mathcal{O}_{\mathbb{P}^{2}}(2))$ $C_{2}+E_{5}$
(I) Conic bundles. In characteristic zero, the discriminant divisor $\Delta_{f}$ of a threefold conic bundle $f:X\to Y$ is reduced and normal crossing. However, both the conclusions fail in characteristic two. Because of this, we shall need some minor modifications of the proofs. For example, if $\rho(X)=3$, $Y=\mathbb{P}^{1}\times\mathbb{P}^{1}$, and $\Delta_{f}\sim\mathcal{O}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{P}^{1}\times\mathbb{P}^{1}}}(2,0)$, then we immediately get a contradiction in characteristic zero, because $\Delta_{f}$ does not contain any smooth rational curve as a connected component [MM83, Proposition 6.2(3) and Proposition 6.3(1)].
@@
Remark 1.3.
The authors do not know whether all classes of Fano threefolds listed in Theorem 1.2 actually exist in arbitrary characteristic (e.g., the existence of Fano threefolds of No. 2-8 is not clear when $p=2$).
Acknowledgements: This paper is based on the master thesis [Asa] of the first author, in which primitive Fano threefolds of odd characteristic are classified. The authors thank the referee for reading the manuscript carefully and for suggesting several improvements. The second author was funded by JSPS KAKENHI Grant numbers JP22H01112 and JP23K03028.
@@ 2.2. Split double covers
    If $p\neq 2$, then the branch divisor $D$ satisfies $\mathcal{O}_{Y}(D)\simeq\mathcal{L}^{\otimes 2}$.
@@
Lemma 2.4.
Let $f:X\to Y$ be a double cover of smooth projective varieties. Set $\mathcal{L}:=(f_{*}\mathcal{O}_{X}/\mathcal{O}_{Y})^{-1}$, which is an invertible sheaf (Remark 2.2). Assume that $p=2$ and $f$ is separable. Then $H^{0}(Y,\mathcal{L})\neq 0$.
Proof.
Note that the trace map ${\operatorname{Tr}}:f_{*}\mathcal{O}_{X}\to\mathcal{O}_{Y}$ is nonzero, because the field-theoretic trace map ${\operatorname{Tr}}:K(X)\to K(Y)$ is given by ${\operatorname{Tr}}(b)=b+\sigma(b)$ for the Galois involution $K(X)\to K(X)$, and hence ${\operatorname{Tr}}(b)\neq 0$ for $b\in K(X)\setminus K(Y)$. As the composite $\mathcal{O}_{Y}$-module homomorphism $\mathcal{O}_{Y}\hookrightarrow f_{*}\mathcal{O}_{X}\xrightarrow{{\operatorname{Tr}}}\mathcal{O}_{Y}$ is zero, we get a nonzero $\mathcal{O}_{Y}$-module homomorphism $\mathcal{L}^{-1}\simeq f_{*}\mathcal{O}_{X}/\mathcal{O}_{Y}\to\mathcal{O}_{Y}$. Therefore, $H^{0}(Y,\mathcal{L})\neq 0$. ∎
@@
    $f$ is separable and $H^{1}(Y,\mathcal{O}_{Y}(E))=0$ for every effective Cartier divisor $E$ on $Y$.
  4. (4)
    $f$ is inseparable and $Y$ is $F$-split, i.e., $\mathcal{O}_{Y}\to F_{*}\mathcal{O}_{Y}$ splits as an $\mathcal{O}_{Y}$-module homomorphism.
@@
It is well known that $f$ is split when (1) holds. In what follows, we assume $p=2$. Note that $\mathcal{O}_{Y}\to f_{*}\mathcal{O}_{X}$ splits if and only if
@@
splits. Since the extension class corresponding to this exact sequence is contained in $H^{1}(Y,\mathcal{L})$, (2) implies that $f$ is split. Note that (3) implies (2) by Lemma 2.4.
Assume (4). Then the absolute Frobenius morphism $F:Y\to Y$ of $Y$ factors through $f:X\to Y$:
@@
and hence the splitting of $\mathcal{O}_{Y}\to F_{*}\mathcal{O}_{Y}$ implies the splitting of $\mathcal{O}_{Y}\to f_{*}\mathcal{O}_{X}$. Therefore, (4) implies that $f$ is split. ∎
@@ 2.3. Brauer groups
Let $f:X\to Y$ be a projective morphism of smooth varieties. Fix $r\in\mathbb{Z}_{>0}$. Assume that ${\operatorname{Br}}(Y)=0$ and every fibre of $f$ is isomorphic to $\mathbb{P}^{r}$. Then there exists a vector bundle $E$ of rank $r+1$ such that $X$ is isomorphic to $\mathbb{P}(E)$ over $Y$.
@@
Note that $f$ is flat [Mat86, Theorem 23.1]. Fix a closed point $y\in Y$ and set $A:=\mathcal{O}_{Y,y}$. For the completion $\widehat{A}=\widehat{\mathcal{O}}_{Y,y}$, we have an isomorphism $X\times_{Y}{\operatorname{Spec}}\,\widehat{A}\simeq\mathbb{P}^{r}_{\widehat{A}}$ over $\widehat{A}$ [Ser06, Corollary 1.2.15]. By Artin’s approximation theorem, we obtain an isomorphism $X\times_{Y}{\operatorname{Spec}}\,A^{h}\simeq\mathbb{P}^{r}_{A^{h}}$ over the henselisation $A^{h}=\mathcal{O}_{Y,y}^{h}$. Therefore, there exists an étale surjective morphism $Y^{\prime}\to Y$ such that $X\times_{Y}Y^{\prime}\simeq\mathbb{P}^{r}_{Y^{\prime}}$.
@@
This implies that the ${\rm PGL}_{r+1}$-torsor $f:X\to Y$ comes from ${\rm GL}_{r+1}$-torsor. Since a ${\rm GL}_{r+1}$-torsor is nothing but a projective space bundle $\mathbb{P}_{Y}(E)$, the assertion holds by the following facts.
@@
  • $H^{1}_{{\operatorname{\acute{e}t}}}(Y,{\rm PGL}_{r+1})$ consists of the isomorphism classes of $\pi:W\to Y$ such that $W$ is a flat proper morphism such that $W\times_{Y}Y^{\prime}\simeq\mathbb{P}^{r}_{Y^{\prime}}$ for some étale surjective morphism $Y^{\prime}\to Y$.
@@ 3.1. Types of extremal rays
    For every curve $C$ on $X$, $[C]\in R$ if and only if $f(C)$ is a point.
@@
    Assume that $R$ is of type $C$. Then the following hold, because $f^{-1}(y)$ is a conic on $\mathbb{P}^{2}$ and we have $-K_{X}\cdot f^{-1}(y)=2$ for every closed point $y\in Y$.
@@ 3.2. Del Pezzo fibrations
    For a closed point $y\in Y$, the effective Cartier divisor $f^{*}y$ is a prime divisor, i.e., every scheme-theoretic fibre of $f$ is geometrically integral.
@@
    $X_{\overline{K}}$ is a projective canonical del Pezzo surface, i.e., $X_{\overline{K}}$ is a projective normal surface such that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X}_{\overline{K}}$ has at worst canonical singularities and $-K_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X}_{\overline{K}}}$ is ample.
@@
    ${\operatorname{Pic}}X_{\overline{K}}$ is a finitely generated free abelian group. Furthermore, ${\operatorname{Pic}}X_{K}\simeq\mathbb{Z}$.
@@
    If $K_{X_{K}}^{2}=9$, then $f$ is a $\mathbb{P}^{2}$-bundle, i.e., there exists a locally free $F$ on $Y$ of rank $3$ such that $X$ is isomorphic to $\mathbb{P}_{Y}(F)$ over $Y$. In particular, every fibre is $\mathbb{P}^{2}$.
@@
$\Delta(X_{\overline{K}},\mathcal{O}_{X}(D)|_{X_{\overline{K}}})=2+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\mathcal{O}_{X}(D)|_{X_{\overline{K}}})}^{2}-h^{0}(X_{\overline{K}},\mathcal{O}_{X}(D)|_{X_{\overline{K}}})=0.$
@@ 3.3. Contraction morphisms
The assertions (1) and (2) follow from [TanI, Theorem 2.4] (cf. [Kaw21, Corollary 3.7]). The assertion (3) holds by $\chi(X,\mathcal{O}_{X})=1$ and $\chi(X,\mathcal{O}_{X})=\frac{1}{24}(-K_{X}\cdot c_{2}(X))$, where the latter one follows from the Riemann–Roch theorem.
@@
On the other hand, we have
@@
    If $X$ is a Fano threefold, then $R^{i}f_{*}\mathcal{O}_{X}=0$ for every $i>0$. Furthermore, $H^{j}(X,\mathcal{O}_{X})\simeq H^{j}(Y,\mathcal{O}_{Y})$ for every $j\geq 0$.
@@
(1) If $\dim X_{y}\leq 1$ for every closed point $y\in Y$, then $R^{i}f_{*}\mathcal{O}_{X}=0$ holds by [Tan15, Theorem 0.5]. We may assume that $\dim X_{y}\geq 2$ for some closed point $y\in Y$. In particular, $\dim Y=0$ or $\dim Y=3$, as we assume $\dim Y\neq 1$. If $\dim Y=0$, then we obtain $R^{i}f_{*}\mathcal{O}_{X}=0$ by Lemma 3.8.
Assume that $\dim Y=3$. By $\dim X_{y}\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}$, $f(E)$ is a point for $E:={\operatorname{Ex}}(f)$. Then we have an exact sequence for every $r\in\mathbb{Z}_{\geq 0}$:
@@
(2) By (1), we may assume that $\dim Y=1$. Pick general closed points $Q_{1},...,Q_{n}$ on $Y$. Set $S_{1}:=f^{*}Q_{1},...,S_{n}:=f^{*}Q_{n}$, and $D_{Y}:=Q_{1}+\cdots+Q_{n}$. By the Serre vanishing theorem
@@
is exact. Since each $S_{i}$ is a canonical del Pezzo surface (Proposition 3.5(2)), we have $H^{i}(S_{\ell},\mathcal{O}_{S_{\ell}})=0$ for every $\ell$ and every $i>0$. It holds that $H^{i}(X,\mathcal{O}_{X})=0$ for every $i>0$, which implies $H^{i}(X,\mathcal{O}_{X}(f^{*}D_{Y}))=0$ for every $i>0$. ∎
@@
Set $\mathcal{O}_{D}(D):=\mathcal{O}_{X}(D)|_{D}$. We have